A ball of mass 8kg falls from rest from a height of 100m. Neglecting air resistance, calculate its kinetic energy after falling a distance of 30m .(Take g as 10ms^-2)


1
Expert's answer
2021-02-18T18:53:20-0500
"KE=\Delta PE=mg(h_2-h_1),""KE=8\ kg\cdot10\ \dfrac{m}{s^2}\cdot(100\ m-70\ m)=2400\ J."